# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ V-C1 Violation of User Privacy
Exposure of Personal Information: A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [292, 172, 273]. In more severe cases, this can extend to financial data like credit card numbers or passwords [12], leading to direct financial loss. What makes this threat particularly potent is the agent’s role as a central data aggregator. Since agents often integrate with multiple user data silos (e.g., email, calendar, cloud storage, and social media), a breach does not just expose isolated pieces of information. Instead, it allows for the exfiltration of a comprehensively aggregated user profile, where the potential for harm far exceeds the sum of its parts.
@@
Behavioral and Psychological Profiling: A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will. Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [63, 91, 278, 274]. These disclosed profiles put users at the risk of manipulation, targeted scams, or social engineering.
V-C2 Psychological and Social Manipulation
@@
Beyond simple data theft, a compromised agent can become a powerful tool for psychological manipulation, exploiting the user’s trust and the agent’s persuasive capabilities. This form of attack targets the user’s beliefs, decisions, and relationships.
@@
Belief and Opinion Shaping: The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time. By personalizing the misinformation to the user’s psychological profile, the agent can effectively manipulate their worldview, influence their voting behavior, or radicalize their beliefs. This exploits the inherent persuasiveness of conversational AI. Park et al. [220] highlights how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect. They note that AI can “simulate empathy” to build rapport before manipulating the user.
@@
Sophisticated Social Engineering and Impersonation: A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.). It can leverage this to conduct highly convincing impersonation attacks. For example, it could send a fraudulent email to the user’s colleague or family member, perfectly mimicking the user’s tone, to request a password reset, a fund transfer, or sensitive information. This attack is far more credible than generic phishing attempts. Greshake et al. [92] demonstrates how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf. This mechanism could be used to weaponize an agent for impersonation.
V-C3 Execution of Malicious and Harmful Tasks
@@
Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [192].
@@
Economic Manipulation: The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts. For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [221]. The harm is often latent and difficult to detect, potentially leading to professional failure or corporate espionage. This extends to broader market manipulation, where an agent could use the user’s social media accounts to automate large-scale disinformation campaigns, such as posting fake product reviews or spreading rumors to affect a company’s stock price, making the user an unwitting accomplice in a larger economic attack.
@@
Malicious Guidance: A compromised agent can also be used as a direct vector for attacking the user’s digital environment. It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [273, 158]. In a more severe scenario, a jailbroken or manipulated agent can bypass its safety protocols to provide overtly harmful instructions. This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [31, 183, 294, 174].
V-D Defense Countermeasure Prospect
We will investigate the possible defense measures that can be taken to address the security risks in the user-agent interaction. Here we only present the defenses for attacks from malicious users; those for compromised agents (against users, other agents, and environments) are uniformly put in Section VI-C.
V-D1 Countermeasures for Text-Based Attacks
@@
To mitigate prompt-based attack risks in user–agent interactions, we hope developers to adopt a multi-layered defense framework targeting three key stages: input/output filtering, external data source evaluation, and internal message isolation.
@@
Input and Output Filtering. Before user inputs are processed by the agent system, multiple approaches can be conducted for a semantic-level input safety review. For example, methods based on intent analysis [349, 293], perplexity calculation [126], and fine-tuned safety classifiers [166, 352, 333, 122] can be employed to identify attack instructions and malicious intentions in the input stage. After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [352, 333, 122, 202, 324], to ensure alignment with safety objectives.
@@
External Source Evaluation. To counter indirect prompt injection attack, external sources (e.g., retrieved documents, web content) should be assessed for safety and trustworthiness [359]. The strategies that can be adopted include: (1) whitelisting verified external sources to block the injection of malicious content; (2) tagging retrieved results with source metadata and risk scores to guide the system to handle potential high-risk content with caution; and (3) sandboxing potential high-risk content to prevent it from entering the model context and affecting the model behavior.
To ensure the effectiveness and comprehensiveness of the aforementioned mechanism in real-world deployment, systems should undergo continuous security evaluation. Boisvert et al. [28] propose DoomArena, an attack-generation framework to test agents against evolving security risks such as prompt injection attacks, helping to uncover vulnerabilities and strengthen defenses against evolving prompt injection threats.
V-D2 Countermeasures for Multimodal Attacks
@@
To address the serious challenges posed by multimodal attacks, relying solely on output-side text-based safety mechanisms is far from sufficient. Future security frameworks must incorporate cross-modal perception and collaborative defense capabilities to effectively detect and intercept malicious attacks launched through non-textual channels. In the following, we explore core defense strategies against multimodal attacks from several key perspectives.
@@
Image Purification: To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals. These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [120, 51, 310]. Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate. In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [210].
@@
Audio Purification: To defend against attacks targeting the audio channel, signal processing techniques can also be applied [226]. Methods such as resampling, injecting slight background noise, altering pitch, or changing playback speed can disrupt the effectiveness of adversarial waveforms, causing them to either fail in automatic speech recognition (ASR) systems or decode into benign content. Moreover, applying band-pass or low-pass filters can eliminate abnormal signals outside the typical human voice frequency range, which are often exploited to carry adversarial perturbations.
@@
Cross-Modal Consistency Verification. The core idea of this defense strategy is to verify whether there is a semantic or intentional conflict between inputs from different modalities. A lightweight, independent cross-modal semantic alignment detection model can be employed [227, 224]. This model takes the embedding vectors of textual prompts and image/audio inputs and determines whether they are semantically aligned. Additionally, before processing user requests, the system can utilize a dedicated vision or audio captioning model to generate a textual description of non-textual inputs. The generated description is then combined with the original user prompt to perform a comprehensive safety evaluation. To counter attacks based on visual text overlays, the system may first run an OCR engine on the image to extract any embedded text. This extracted text can be merged with the user’s original prompt and passed through a text-based safety filter. This approach effectively converts risks from non-textual modalities into the textual domain, allowing mature text safety techniques to be leveraged for defense.
V-D3 Countermeasures for Privacy Leakage
@@
To address the privacy leakage risks that arise in user-agent interaction, we propose the following privacy protection defense strategies.
@@
Data Minimization and Anonymization. During the multimodal data collection phase, a strict data minimization principle should be enforced, ensuring that only the information necessary for task completion is collected. Sensitive biometric data (e.g., facial features, voiceprints, gesture patterns) should be processed using differential privacy or k-anonymity techniques to mitigate the risk of identity reconstruction. Besides, a hierarchical data access control mechanism should be established to ensure that each system component can access only the minimal dataset required for its functionality. To protect sensitive biometric features such as facial information, Wen et al. [301] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
@@
Privacy Leakage Prompt detection. A multi-layered input validation and filtering mechanism based on semantic analysis and intent recognition should be established to detect and block adversarial prompts that attempt to induce the system to leak sensitive information. For example, the GenTel-Shield [166] defense module incorporates semantic feature extraction and intent classification to identify potential privacy leakage attacks within user inputs. Evaluated on the large-scale benchmark dataset GenTel-Bench, GenTel-Shield demonstrates strong detection performance and represents one of the most practical and effective solutions in this domain.
@@
V-D4 Countermeasures For DoS
To address the Denial of Service risks in user-agent interaction, we propose the following privacy protection defense strategies.
Resource Management and Anomaly Detection. Fine-grained resource quota management should be implemented by setting computation limits for each user session and agent instance. An output length prediction algorithm can be introduced to monitor and truncate potentially malicious long outputs in real time during the generation process. In addition, a real-time monitoring mechanism should be established to track request frequency and resource consumption from individual users or IP addresses, enabling adaptive adjustments to model responses or temporary access restrictions for suspicious users.
Efficient Reasoning Compression. To defend against OverThink attacks, a promising direction is to improve communication efficiency by reducing the token consumption in the reasoning process. Recent studies have shown that effective reasoning does not necessarily require verbose Chain-of-Thought (CoT) traces to maintain performance. For example, LightThinker [339] proposes a step-by-step compression method that condenses intermediate reasoning into shorter yet semantically equivalent representations, significantly reducing inference cost without compromising accuracy. GoGI-Skip [364] leverages goal-gradient importance signals to dynamically skip low-value reasoning steps, reducing token usage while preserving performance. Compressed CoT (CCoT) [40] introduces variable-length, information-dense “thought tokens” as a compact and controllable alternative to traditional textual reasoning chains. C3oT [135] trains the model on paired long and short CoT examples, enabling it to generate compressed reasoning traces during inference under specific control prompts. Integrating these lightweight reasoning mechanisms into agent communication protocols can significantly enhance inference efficiency, reduce computational latency, and mitigate the resource overhead caused by adversarial bait tasks. Furthermore, applying techniques such as dynamic reasoning budget constraints, step skipping, or output summarization during generation can effectively truncate excessively verbose reasoning chains, thereby preserving responsiveness and resource availability under adversarial conditions. These strategies not only improve system robustness against slowdown attacks but also enhance overall communication efficiency in both multi-agent and human-agent interactions.
Model Robustness Enhancement. To enhance model robustness, adversarial examples should be incorporated during the training and fine-tuning stages, enabling the model to recognize malicious inputs that may contain Denial-of-Service (DoS) triggers. Furthermore, a behavior-constrained system based on anomaly detection can be deployed during inference, which performs output validity checks to detect repetitive, nonsensical, or abnormally long responses, thereby preventing the model from generating overtly anomalous outputs.
V-E Takeaways
User-agent interaction enables agents to process multimodal inputs, such as text, images, audio, and their combinations. Since this process needs to directly face the uncertain inputs from diverse users, its security risks are severe. We divide the existing risks into risks from malicious users and risks from compromised agents. Then we detailedly discuss the defense countermeasures for risks from malicious users. Defenses for compromised agents will be introduced in the agent-agent communication section. Overall, user-agent interaction is critical for bridging human intent with agent execution, but its security requires long-term study.

### v4

@@ V-B3 Risks from Semantic Interpretation Layer (L3)
Exposure of Personal Information. A compromised agent can be induced to leak the user’s Personally Identifiable Information (PII) that it has access to, such as the user’s name, email, address, and conversation history [wang2025pig, liao2024eia, triedman2025multi]. In more severe cases, this can extend to financial data like credit card numbers or passwords [alizadeh2025simple], leading to direct financial loss. What makes this threat particularly potent is the agent’s role as a central data aggregator. Since agents often integrate with multiple user data silos (e.g., email, calendar, cloud storage, and social media), a breach does not just expose isolated pieces of information. Instead, it allows for the exfiltration of a comprehensively aggregated user profile, where the potential for harm far exceeds the sum of its parts.
@@
Behavioral and Psychological Profiling. A compromised agent can be manipulated to analyze the user’s inputs across sessions to build a detailed behavioral or psychological profile against their will. Moreover, a more insidious risk happens where the agent deduces highly sensitive attributes (e.g., health conditions, political affiliations, or undisclosed personal relationships) from seemingly innocuous conversational data that the user never explicitly provided [du2025automated, green2025leaky, wang2025unveiling, tshimula2024psychological]. These disclosed profiles put users at risk of manipulation, targeted scams, or social engineering.
@@
R4: Psychological and Social Manipulation. Beyond simple data theft, a compromised agent can become a powerful tool for psychological manipulation, exploiting the user’s trust and the agent’s persuasive capabilities. This form of attack targets the user’s beliefs, decisions, and relationships.
@@
Belief and Opinion Shaping. The agent can be instructed to subtly introduce biased information, conspiracy theories, or political propaganda into its responses over time. By personalizing the misinformation to the user’s psychological profile, the agent can effectively manipulate their worldview, influence their voting behavior, or radicalize their beliefs. This exploits the inherent persuasiveness of conversational AI. Park et al. [park2024ai] highlight how models can be used for manipulative purposes, including generating persuasive, deceptive content that is difficult for humans to detect. They note that AI can “simulate empathy” to build rapport before manipulating the user.
@@
Sophisticated Social Engineering and Impersonation. A compromised agent has intimate knowledge of the user’s communication style, vocabulary, and relationships (gleaned from emails, messages, etc.). It can leverage this to conduct highly convincing impersonation attacks. For example, it could send a fraudulent email to the user’s colleague or family member, perfectly mimicking the user’s tone, to request a password reset, a fund transfer, or sensitive information. This attack is far more credible than generic phishing attempts. Greshake et al. [greshake2023not] demonstrate how an agent can be poisoned by external data (like a webpage it summarizes) and then turned against the user or even used to attack other systems on the user’s behalf. This mechanism could be used to weaponize an agent for impersonation.
@@
R5: Execution of Malicious and Harmful Tasks. Once compromised, an agent can be weaponized, transforming from a trusted assistant into an active executor of malicious tasks that can sabotage the user’s interests or directly endanger them, representing a significant risk escalation [mai2025you].
@@
Economic Manipulation. The agent can be instructed to inflict subtle yet significant damage in professional or economic contexts. For a user relying on it for work, it could covertly introduce logical errors into computer code, provide flawed data in financial projections, or leak confidential business strategies discussed in conversations [pearce2025asleep]. The harm is often latent and difficult to detect, potentially leading to professional failure or corporate espionage. This extends to broader market manipulation, where an agent could use the user’s social media accounts to automate large-scale disinformation campaigns, such as posting fake product reviews or spreading rumors to affect a company’s stock price, making the user an unwitting accomplice in a larger economic attack.
@@
Malicious Guidance. A compromised agent can also be used as a direct vector for attacking the user’s digital environment. It can be triggered to generate scripts that download malware, trick the user into visiting phishing websites, or send highly convincing phishing emails on the user’s behalf, thereby damaging their reputation and spreading the attack to their contacts [triedman2025multi, lee2024prompt]. In a more severe scenario, a jailbroken or manipulated agent can bypass its safety protocols to provide overtly harmful instructions. This includes generating tutorials for synthesizing toxic substances, creating malicious code on demand, or providing dangerously flawed medical or financial advice, directly threatening the user’s physical safety and financial stability [chao2023jailbreaking, liu2023jailbreaking, wang2024jailbreak, lin2024figure].
V-C Defense Countermeasure Prospect
We investigate the possible defense measures that can address the security risks in the user-agent interaction.
V-C1 Defenses on Data Transmission Layer (L1)
To mitigate the risks on L1, developers should focus on the following aspects.
D1: Encryption and Traffic Obfuscation. Data transmission needs to strictly adopt the HTTPS protocol when building communication channels, perform end-to-end encryption on user input data and agent response content. At the same time, it is necessary to introduce a traffic obfuscation mechanism. By dynamically adjusting the length of data packets and adding random padding fields, it can resist analytical attacks based on traffic fingerprints and cut off the possibility of attackers inferring interaction content or system architecture through traffic patterns.
D2: Integrity Authentication. Deploying digital signature mechanisms based on asymmetric encryption algorithms is also necessary. The generated signature can avoid data tampering caused by MITM attacks and verify the legitimacy of the sender’s public key. At the same time, a dual verification mechanism of timestamps and random numbers should be introduced. The timestamp is to ensure data timeliness, and the random numbers can resist malicious interference of replay attacks on interaction sequences.
D3: DoS Mitigation. Traffic filtering systems are needed on the server side. For example, a normal traffic AI model automatically filters abnormal traffic such as DDoS attacks and malicious crawlers. Meanwhile, a dynamic rate-limiting strategy is set up to allocate differentiated traffic quotas based on dimensions like user level and interaction frequency, preventing malicious requests from occupying all resources. At the client level, strict thresholds for the size of received data and format verification rules are established. Oversized payloads that exceed the threshold are automatically truncated, while malformed data that does not conform to the preset format is directly rejected. In addition, a multi-node redundant communication architecture can also work, which automatically switches to a standby node when the main communication node is flooded, ensuring service continuity.
V-C2 Defenses on Interaction Protocol Layer (L2)
To mitigate the risks from L2, defenses must be deeply integrated into the protocol’s design, focusing on robust authentication, strict state integrity, and zero-trust authorization.
D1: Identity Authentication. To defend against identity spoofing, the interaction protocol must employ multi-factor, continuous, and context-aware authentication rather than relying solely on static credentials. Strong MFA (e.g., cryptographic tokens or device-bound certificates) can prevent attackers with stolen credentials from initiating malicious sessions, while continuous authentication mechanisms (e.g., behavioral profiling or session-binding tokens) ensure that the session remains tied to the legitimate user throughout its lifetime. Additionally, any anomalous patterns should trigger real-time verification, effectively limiting the attacker’s ability to hijack the agent even when initial credentials have been compromised.
D2: Session Status Verification. Preventing session manipulation requires strict server-side enforcement of protocol state transitions, ensuring that each event in the interaction stream is both syntactically valid and semantically consistent with the session’s current state. The protocol should incorporate explicit state-transition automata, sequence numbers, and causal-dependency checks so that logically inconsistent event sequences can be rejected before reaching the agent. Additionally, integrity mechanisms such as event-stream hashing, tamper-evident logs, and runtime invariants can guarantee that the client and server states remain synchronized, eliminating opportunities for attackers to desynchronize the protocol logic or bypass safety checks.
D3: Authority Verification. Defending against authority abuse requires enforcing strict context isolation and privilege verification when session context is propagated across agents. Each receiving agent must re-validate the authorization scope independently rather than inheriting trust from upstream agents. The protocol should also mandate that context-passing events include cryptographically verifiable provenance and explicit privilege boundaries, preventing attackers from crafting manipulated contexts that appear legitimate. By ensuring that privilege escalation cannot occur through implicit trust between agents, the system can maintain robust security boundaries in multi-agent workflows.
V-C3 Defenses on Semantic Interpretation Layer (L3)
The defenses on L3 are classified based on the risks.
(i) Against Risks from Malicious Users.
@@ V-C3 Defenses on Semantic Interpretation Layer (L3)
D1: Defense for Text-based Attacks To mitigate text-based attack risks in user-agent interactions, developers should adopt a multi-layered defense framework targeting three key stages: input/output filtering, external data source evaluation, internal metadata isolation, and cognitive load regulation.
@@
Input and Output Filtering. Before user inputs are processed by the agent system, multiple approaches can be conducted for a semantic-level input safety review. For example, methods based on intent analysis [zhang2024intention, wang2024defending], perplexity calculation [jain2023baseline], and fine-tuned safety classifiers [li2024gentel, zhang2024shieldlm, zeng2025shieldgemma2robusttractable, inan2023llama] can be employed to identify attack instructions and malicious intentions in the input stage. After generating the final response, it is also necessary to go through an output review mechanism, such as specific output safety detection models [zhang2024shieldlm, zeng2025shieldgemma2robusttractable, inan2023llama, mazeika2024harmbench, yu2024gptfuzzer], to ensure alignment with safety objectives.
@@
External Source Evaluation. To counter indirect prompt injection attack, external sources (e.g., retrieved documents, web content) should be assessed for safety and trustworthiness [zhou2025trustrag]. The strategies that can be adopted include: (1) whitelisting verified external sources to block the injection of malicious content; (2) tagging retrieved results with source metadata and risk scores to guide the system to handle potential high-risk content with caution; and (3) sandboxing potential high-risk content to prevent it from entering the model context and affecting the model behavior.
Internal Metadata Isolation. Since user-to-agent privacy leakage attacks generally exploit prompt-based information extraction, their defense strategies strongly parallel those for prompt injection. The system should restrict internal metadata exposure, ensuring that system prompts, roles, tool configurations, and agent topologies are never treated as answerable content. Output-level safety filtering further blocks inadvertent disclosure, helping detect adversarial attempts that induce the model to reveal private information.
Cognitive Load Regulation. To defend against exhaustion attacks, the system should implement fine-grained resource quotas, dynamic token budgets, and real-time output-length prediction to detect and truncate malicious long-form responses. For attacks like OverThink, lightweight reasoning methods—such as compressed CoT, selective reasoning, or adaptive reasoning depth control—should be integrated to suppress unnecessary cognitive loops. Runtime behavioral detectors should halt repetitive, nonsensical, or computationally excessive outputs, ensuring service availability and preventing adversarial prompt-induced degradation of inference performance.
@@
D2: Defense for Multimodal Attacks. To address multimodal attacks, future security frameworks must incorporate cross-modal perception and collaborative defense capabilities to effectively detect and intercept malicious attacks launched through non-textual channels. In the following, we explore core defense strategies against multimodal attacks from three key perspectives.
@@
Image Purification. To counter visual perturbations and camouflage-based attacks, various image processing techniques can be employed to disrupt or eliminate adversarial signals. These include simple transformations such as random resizing, cropping, rotation, or mild JPEG compression [ilyas2019adversarial, das2017keeping, xie2017mitigating]. Although lightweight, such operations can significantly degrade pixel-level adversarial patterns meticulously crafted by attackers, thereby reducing the attack success rate. In addition, diffusion models can be used to reconstruct the input image, effectively “washing out” subtle and imperceptible adversarial perturbations [nie2022diffusion].
@@
Audio Purification. To defend against attacks targeting the audio channel, signal processing techniques can also be applied [qin2019imperceptible]. Methods such as resampling, injecting slight background noise, altering pitch, or changing playback speed can disrupt the effectiveness of adversarial waveforms, causing them to either fail in automatic speech recognition (ASR) systems or decode into benign content. Moreover, applying band-pass or low-pass filters can eliminate abnormal signals outside the typical human voice frequency range, which are often exploited to carry adversarial perturbations.
@@
Cross-Modal Consistency Verification. The core idea of this defense strategy is to verify whether there is a semantic or intentional conflict between inputs from different modalities. A lightweight, independent cross-modal semantic alignment detection model can be employed [radford2021learning, poppi2024safe]. This model takes the embedding vectors of textual prompts and image/audio inputs and determines whether they are semantically aligned. Additionally, before processing user requests, the system can utilize a dedicated vision or audio captioning model to generate a textual description of non-textual inputs. The generated description is then combined with the original user prompt to perform a comprehensive safety evaluation. To counter attacks based on visual text overlays, the system may first run an OCR engine on the image to extract any embedded text. This extracted text can be merged with the user’s original prompt and passed through a text-based safety filter. This approach effectively converts risks from non-textual modalities into the textual domain, allowing mature text safety techniques to be leveraged for defense.
(ii) Against Risks from Malicious Agents.
@@
D3: Defense for Privacy Violation. To address the privacy leakage risks that arise in user-agent interaction, we propose the following privacy protection defense strategies.
@@
Data Minimization and Anonymization. During the multimodal data collection phase, a strict data minimization principle should be enforced, ensuring that only the information necessary for task completion is collected. Sensitive biometric data (e.g., facial features, voiceprints, gesture patterns) should be processed using differential privacy or k-anonymity techniques to mitigate the risk of identity reconstruction. Besides, a hierarchical data access control mechanism should be established to ensure that each system component can access only the minimal dataset required for its functionality. To protect sensitive biometric features such as facial information, Wen et al. [wen2022identitydp] propose a differential privacy-based anonymization framework, IdentityDP, to effectively safeguard identity information while preserving visual utility and task performance, offering a practical solution for privacy protection in multimodal systems.
@@
Privacy Leakage Prompt detection. A multi-layered input validation and filtering mechanism based on semantic analysis and intent recognition should be established to detect and block adversarial prompts that attempt to induce the system to leak sensitive information. For example, the GenTel-Shield [li2024gentel] defense module incorporates semantic feature extraction and intent classification to identify potential privacy leakage attacks within user inputs. Evaluated on the large-scale benchmark dataset GenTel-Bench, GenTel-Shield demonstrates strong detection performance and represents one of the most practical and effective solutions in this domain.
@@
D4: Behavior Auditing and Restrictions. To counter manipulation risks in agent-to-user interactions, the system should adopt a safety framework that limits persuasive personalization, constrains user modeling, and continuously audits agent behavior. The agent must be prevented from leveraging psychological traits or roleplay profiling to generate tailored content, particularly in politically, ideologically, or emotionally sensitive domains. Safety classifiers and style-consistency filters can detect manipulative strategies that aim to influence the user’s worldview or decisions. These integrated controls prevent a compromised agent from exploiting user trust, shaping cognition, or weaponizing knowledge of the user’s communication patterns.
D5: Capability Boundary Control. Mitigating the threat of a weaponized agent requires strict capability scoping, controlled tool usage, and rigorous verification of high-risk actions. The system should enforce fine-grained capability boundaries to ensure that the agent cannot autonomously initiate sensitive operations without explicit, authenticated user approval. At the same time, tool-use governance must incorporate permission checks and parameter constraints to prevent covert sabotage in high-stakes domains. In addition, safety filters should strictly block any form of malicious or high-risk content—including malware, phishing attempts, disinformation, or dangerous procedural instructions—regardless of how the user phrases the request, to ensure reliability and prevent harmful misuse.
V-D Takeaways
The security of user-agent interaction highly relies on the collaborative defenses of the three-layered architecture. As the foundational communication layer, L1 needs to resist risks such as eavesdropping and traffic analysis through HTTPS encryption and traffic obfuscation. On L2, developers should introduce identity and session security mechanisms in the protocol design, such as authentication and session state verification, to prevent identity spoofing and privilege abuse; L3 is the high-risk layer. As a result, it requires defenses such as input semantic filtering and cross-modal consistency verification. This layered architecture not only clarifies the location/reason of each risk but also provides a technical framework for precise defense.